# https://ldtho.github.io/DifFUSER/

### v1

@@ 5 Dicussion and Comparison with State-of-the-arts
Methods Mod.normal-$\uparrow$ NDSnormal-$\uparrow$ mAPnormal-$\uparrow$ mATEnormal-$\downarrow$ mASEnormal-$\downarrow$ mAOEnormal-$\downarrow$ mAVEnormal-$\downarrow$ mAAEnormal-$\downarrow$ PointPainting [38] LC 61.0 54.1 0.380 0.260 0.541 0.293 0.131 PointAugmenting [39] LC 71.1 66.8 0.253 0.235 0.354 0.266 0.123 MVP [49] LC 70.5 66.4 0.263 0.238 0.321 0.313 0.134 FusionPainting [44] LC 71.6 68.1 0.256 0.236 0.346 0.274 0.132 UVTR [13] LC 71.1 67.1 0.306 0.245 0.351 0.225 0.124 TransFusion [1] LC 71.7 68.9 0.259 0.243 0.359 0.288 0.127 BEVFusion [16] LC 71.8 69.2 0.260 0.250 0.370 0.270 0.130 BEVFusion [21] LC 72.9 70.2 0.261 0.239 0.329 0.260 0.134 DeepInteraction [47] LC 73.4 70.8 0.257 0.240 0.325 0.245 0.128 LinK [25] LC 73.4 69.8 0.258 0.229 0.312 0.234 0.136 LargeKernel-F [6] LC 74.1 71.1 0.236 0.228 0.298 0.241 0.131 CMT [45] LC 74.1 72.0 0.279 0.235 0.308 0.259 0.112 DifFUSER (ours) LC 73.8 71.3 0.250 0.241 0.317 0.244 0.130 Improvement $\uparrow$1.9 $\uparrow$1.0 $\downarrow$0.011 $\uparrow$0.002 $\downarrow$0.012 $\downarrow$0.016 $\downarrow$0.004
Table 3: Results for 3D object detection on NuScenes (val). Bold, blue, and underlined values indicate the best, second-best, and baseline performance, respectively. Mod. stands for Modality. Methods Mod. NDSnormal-$\uparrow$ mAPnormal-$\uparrow$ TransFusion [1] L 70.1 65.1 FUTR3D [4] LC 68.3 64.5 UVTR [13] LC 70.2 65.4 TransFusion [1] LC 71.3 67.5 BEVFusion [21] LC 71.4 68.5 DeepInteration [47] LC 72.6 69.9 CMT [45] LC 72.9 70.3 DifFUSER LC 72.6 70.0 Improvement $\uparrow$1.2 $\uparrow$1.8
Figure 5: Performance Comparison of DifFUSER with and without Sensor Drop (either Lidar or Camera) [Uncaptioned image]
@@ 6 Ablation Study
Table 4: Performance comparison of different sampling methods at varying steps Sampling step 1 2 4 8 DDIM[35] 63.88 66.65 68.82 69.05 DPM-Solver[24] 64.02 66.20 68.90 68.95 DeIS[50] 63.80 66.30 68.81 69.00
Table 5: Multi-task joint training performance Task NDS mIOU Detection 69.59 - Segmentation - 65.8 Joint 68.23 61.36
Table 7: GSM diffusion block design.
Table 6: Comparison of Detection and Segmentation Performance Component mAP NDS mIOU Baseline 68.9 71.4 62.7 +cMiniBiFPN 69.4 71.8 66.5 +GSM (4-step) 69.7 72.1 68.82 +GSM (8-step) 70.0 72.6 69.1
Scale Shift Gate mIOU 59.92 60.73 62.17
Table 7: GSM diffusion block design.
GSM diffusion block design. We conducted experiments about the components of the GSM module, where all models are trained with a shorter training schedule (6 epochs). As shown in Tab. 7, by incrementally incorporating Scale, Shift, and Gate operations, we observed a marked improvement in mean Intersection over Union (mIOU), with the Gate operation significantly boosting the mIOU to 62.17. This underscores the effectiveness of GSM’s dynamic feature modulation, especially its capacity for strong conditioning, which selectively emphasises relevant features, enabling the model to generate meaningful synthetic features.
Contribution of each component. We investigate the contribution of each component in DifFUSER to understand whether the improvement comes from the enhanced cMini-BiFPN fusion architecture or the gated self-conditioned latent diffusion module. We conduct experiments on the NuScenes val set with the same settings as in Sec. 4, using single-task models. It can be seen that by adding BiFPN-like feature fusion module, the performance of BEV map segmentation and 3D object detection are improved by 3.8% in mIOU and 0.4% in NDS, respectively. Further integrating the gated self-conditioned latent diffusion module with a 4-step sampling improves mIOU by an additional 1.3% and NDS by an additional 0.3%. Extending the diffusion sampling to 8 steps further raises mIOU to 69.05% and NDS to 72.6%, which is the final performance of DifFUSER.
@@
Multi-task joint training. Finally, we investigate the performance of DifFUSER in multi-task joint training. We use a shorter training schedule and re-scale the losses of each task to balance the training. The results are shown in Sec. 6. It can be seen that the performance of both tasks decreases compared to single-task training, an issue commonly known as negative transfer, which we leave for future investigation.
@@ 7 Conclusion
In this work, we propose DifFUSER, a diffusion-based generative model for multi-modal fusion in 3D perception. DifFUSER is designed to enhance the quality of fused features by improving the fusion architecture and leveraging the denoising property of diffusion models. Central to our approach is the introduction of a Gated Self-Conditioned Modulated (GSM) latent diffusion module and PSDT paradigm, specifically engineered to improve the fusion output features and robustness against sensor noise and failure. Additionally, cMini-BiFPN fusion architecture emerged as promising multi-scale latent diffusion module. Extensive experiments on the NuScenes dataset show that DifFUSER achieves SOTA performance in BEV map segmentation tasks and can fairly compete with SOTA transformer-based fusion methods in 3D object detection. This shows the potential of generative models to significantly enhance the capability of autonomous driving systems. Our approach opens a direction of using diffusion models for better accuracy and reliability in 3D perception tasks, promising further enhancements and broader applicability of diffusion-based models in complex, real-world environments.
@@ 8.1 Formulation of Diffusion Models
Here, we review the formulation of diffusion models used in DiFFUSER, which based on the frameworks established in [8, 35, 9]. Beginning with an original latent features $\bm{x}_{0}\sim q(\bm{x}_{0})$, we apply a sequential noising mechanism $q$, which incrementally infuses Gaussian noise into latent features $\bm{x}_{1}$, $\bm{x}_{2}$, …, $\bm{x}_{T}$ at each time $t$. Specifically, the noise infusion at each stage is determined by a variance schedule $\beta_{t}\in(0,1)$:
$q(\bm{x}_{1:T}|\bm{x}_{0})$ $=\prod^{T}_{t=1}q(\bm{x}_{t}|\bm{x}_{t-1})$ (9)
$q(\bm{x}_{t}|\bm{x}_{t-1})$ $=\mathcal{N}(\bm{x}_{t};\sqrt{1-\beta_{t}}\bm{x}_{t-1},\beta_{t}\bm{I})$ (10)
Following DDPM [8], instead of iteratively apply $q$, it is feasible to sample $\bm{x}_{t}$ at any chosen $t$ directly:
$q(\bm{x}_{t}|\bm{x}_{0})$ $=\mathcal{N}(\bm{x}_{t};\sqrt{\bar{\alpha}_{t}}\bm{x}_{0},(1-\bar{\alpha}_{t})\bm{I})$ (11)
$=\sqrt{\bar{\alpha}_{t}}\bm{x}_{0}+\epsilon\sqrt{1-\bar{\alpha}_{t}},\epsilon\sim\mathcal{N}(0,\bm{I})$ (12)
where $\bar{\alpha}_{t}=\prod_{s=0}^{t}\alpha_{s}$ and $\alpha_{t}=1-\beta_{t}$. In this context, $\bar{\alpha}_{t}$ is employed for scheduling the noise. According to Bayes’ theorem, the posterior $q(\bm{x}_{t-1}|\bm{x}_{t},\bm{x}_{0})$ is also Gaussian:
$q(\bm{x}_{t-1}|\bm{x}_{t},\bm{x}_{0})$ $=\mathcal{N}(\bm{x}_{t-1};\mu_{t}^{\prime}(\bm{x}_{t},\bm{x}_{0}),\beta_{t}^{\prime}\mathbf{I})$ (13)
@@
$\mu_{t}^{\prime}(\bm{x}_{t},\bm{x}_{0})$ $=\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\bm{x}_{0}+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\bm{x}_{t}$ (14)
@@
To generate a sample from $q(\bm{x}_{0})$, we may sample from $q(\bm{x}_{T})$ and iteratively apply the reverse transitions $q(\bm{x}_{t-1}|\bm{x}_{t})$ back to $\bm{x}_{0}$. Given a sufficiently extensive $T$ and a suitably chosen $\beta_{t}$ series ($\beta_{t}\rightarrow 0$), the distribution $q(\bm{x}_{T})$ converges closely to an isotropic Gaussian. Specifically, our conditional diffusion model is parametrized through $p_{\theta}(x_{t-1}|x_{t},t,\tilde{x}_{0})$. Because $p_{\theta}(x_{t-1}|x_{t},t,\tilde{x}_{0})$ is not be easily estimated but could be approximated using a neural network ( i.e. DifFUSER), where the aim is to predict $\mu_{\theta}$ and $\Sigma_{\theta}$:
$p_{\theta}(\bm{x}_{t-1}|x_{t},t,\tilde{x}_{0})$ $\coloneqq\mathcal{N}(\bm{x}_{t-1};\mu_{\theta}(x_{t},t,\tilde{x}_{0}),\Sigma_{\theta}(x_{t},t,\tilde{x}_{0}))$ (16)
Here, a neural network can also be optimized to predict either $\epsilon$ or $\bm{x}_{0}$ [35, 9]. We opt for predicting $\bm{x}_{0}$ in this paper.
@@ 8.2 Pseudo-code of DifFUSER training and sampling
    $x_{0}\leftarrow$ concatenate(cam_feats, lidar_feats)
    diffusion_target $\leftarrow x_{0}$
@@
    $\tilde{x}_{0}\leftarrow\text{PDST}(x_{0})$
@@
    $x_{t}$leftarrow sqrt(alpha_cumprod(t) * $x_{0}$ + sqrt(1-alpha_cumprod(t)) * noise
@@
    $x_{0}^{\text{pred}}$leftarrow DifFUSER$(\tilde{x}_{0},t)$
@@
    pred $\leftarrow$ task_head($x_{0}^{\text{pred}}$)
@@
    loss $\leftarrow$ diffusion_loss(diffusion_target, $x_{0}^{\text{pred}}$)
@@
    $x_{0}\leftarrow$ concatenate(cam_feats, lidar_feats)
@@
        $x_{0}^{pred}\leftarrow$ DifFUSER($x_{0},x_{t},$ t_now)
    # estimate $x_{t}$ at t_next, sampler can be [35], [50], or [24]
        $x_{t}\leftarrow$ sampler_step($x_{0}^{pred}$, t_now, t_next)
@@
    pred $\leftarrow$ task_head($x_{0}^{\text{pred}}$)

### v2

@@ 5 Dicussion and Comparison with State-of-the-arts
[…]
Table 3: Results for 3D object detection on NuScenes (val). Bold, blue, and underlined values indicate the best, second-best, and baseline performance, respectively. Mod. stands for Modality.
[…]
Figure 5: Performance Comparison of DifFUSER with and without Sensor Drop (either Lidar or Camera)
@@ 6 Ablation Study
GSM diffusion block design. We conducted experiments about the components of the GSM module, where all models are trained with a shorter training schedule (6 epochs). As shown in Tab. 7, by incrementally incorporating Scale, Shift, and Gate operations, we observed a marked improvement in mean Intersection over Union (mIOU), with the Gate operation significantly boosting the mIOU to 62.17. This underscores the effectiveness of GSM’s dynamic feature modulation, especially its capacity for strong conditioning, which selectively emphasises relevant features, enabling the model to generate meaningful outputs.
Contribution of each component. We investigate the contribution of each component in DifFUSER to understand whether the improvement comes from the enhanced cMini-BiFPN fusion architecture or the gated self-conditioned latent diffusion module. We conduct experiments on the NuScenes val set with the same settings as in Sec. 4, using single-task models. It can be seen that by adding BiFPN-like feature fusion module, the performance of BEV map segmentation and 3D object detection are improved by 3.8% in mIOU and 0.4% in NDS, respectively. Further integrating the GSM module with a 4-step sampling improves mIOU by an additional 2.3% and NDS by an additional 0.3%. Extending the diffusion sampling to 8 steps further raises mIOU to 70.04% and NDS to 72.6%, which is the final performance of DifFUSER.
@@
Table 4: Performance comparison of different sampling methods at varying steps
Sampling step 1 2 4 8 DDIM[37] 63.88 66.65 68.82 70.04 DPM-Solver[26] 64.02 66.20 68.90 70.02 DeIS[52] 63.80 66.30 68.81 69.95
Table 5: Multi-task joint training performance
Task NDS mIOU Detection 69.59 - Segmentation - 65.8 Joint 68.23 61.36
Table 6: Comparison of Detection and Segmentation Performance
Component mAP NDS mIOU Baseline** 68.9 71.4 62.7 +cMiniBiFPN** 69.4 71.8 66.5 +GSM (4-step) + PSDT* 69.7 72.1 68.82 +GSM (8-step) + PSDT* 70.0 72.6 70.04 [*] With $\mathcal{L}_{diffusion}$, [**] Without $\mathcal{L}_{diffusion}$
Table 7: GSM diffusion block design.
Scale Shift Gate mIOU 59.92 60.73 62.17
@@
Multi-task joint training. Finally, we investigate the performance of DifFUSER in multi-task joint training. We use a shorter training schedule and re-scale the losses of each task to balance the training. As shown in Sec. 6, the performance of both tasks decreases compared to single-task training, an issue commonly known as negative transfer, which we leave for future investigation.
@@ 7 Conclusion
In this work, we propose DifFUSER, a diffusion-based generative model for multi-modal fusion in 3D perception. DifFUSER is designed to enhance the quality of fused features by improving the fusion architecture and leveraging the denoising property of diffusion models. Central to our approach is the introduction of a Gated Self-Conditioned Modulated (GSM) latent diffusion module and PSDT paradigm, specifically engineered to improve the fusion output features and robustness against sensor noise and failure. Additionally, cMini-BiFPN fusion architecture emerged as a promising multi-scale latent diffusion module. Extensive experiments on the NuScenes dataset show that DifFUSER achieves SOTA performance in BEV map segmentation tasks and can fairly compete with SOTA transformer-based fusion methods in 3D object detection. This shows the potential of generative models to significantly enhance the capability of autonomous driving systems. Our approach opens a direction of using diffusion models for better accuracy and reliability in 3D perception tasks, promising further enhancements and broader applicability of diffusion-based models in complex, real-world environments.
Acknowledgments. This work has been partially funded by The Australian Research Council Discovery Project (ARC DP2020102427). We acknowledge the partial sponsorship of our research by DARPA Assured Neuro Symbolic Learning and Reasoning (ANSR) program, under award number FA8750-23-2-1016.
@@ 8.1 Formulation of Diffusion Models
Here, we review the formulation of diffusion models used in DiFFUSER, which based on the frameworks established in [9, 37, 10]. Beginning with an original latent features $\bm{x}^{F}_{0}\sim q(\bm{x}^{F}_{0})$, we apply a sequential noising mechanism $q$, which incrementally infuses Gaussian noise into latent features $\bm{x}^{F}_{1}$, $\bm{x}^{F}_{2}$, …, $\bm{x}^{F}_{T}$ at each time $t$. Specifically, the noise infusion at each stage is determined by a variance schedule $\beta_{t}\in(0,1)$:
$q(\bm{x}^{F}_{1:T}|\bm{x}^{F}_{0})$ $=\prod^{T}_{t=1}q(\bm{x}^{F}_{t}|\bm{x}^{F}_{t-1})$ (9)
$q(\bm{x}^{F}_{t}|\bm{x}^{F}_{t-1})$ $=\mathcal{N}(\bm{x}^{F}_{t};\sqrt{1-\beta_{t}}\bm{x}^{F}_{t-1},\beta_{t}\bm{I})$ (10)
Following DDPM [9], instead of iteratively apply $q$, it is feasible to sample $\bm{x}^{F}_{t}$ at any chosen $t$ directly:
$q(\bm{x}^{F}_{t}|\bm{x}^{F}_{0})$ $=\mathcal{N}(\bm{x}^{F}_{t};\sqrt{\bar{\alpha}_{t}}\bm{x}^{F}_{0},(1-\bar{\alpha}_{t})\bm{I})$ (11)
$=\sqrt{\bar{\alpha}_{t}}\bm{x}^{F}_{0}+\epsilon\sqrt{1-\bar{\alpha}_{t}},\epsilon\sim\mathcal{N}(0,\bm{I})$ (12)
where $\bar{\alpha}_{t}=\prod_{s=0}^{t}\alpha_{s}$ and $\alpha_{t}=1-\beta_{t}$. In this context, $\bar{\alpha}_{t}$ is employed for scheduling the noise. According to Bayes’ theorem, the posterior $q(\bm{x}^{F}_{t-1}|\bm{x}^{F}_{t},\bm{x}^{F}_{0})$ is also Gaussian:
$q(\bm{x}^{F}_{t-1}|\bm{x}^{F}_{t},\bm{x}^{F}_{0})$ $=\mathcal{N}(\bm{x}^{F}_{t-1};\mu_{t}^{\prime}(\bm{x}^{F}_{t},\bm{x}^{F}_{0}),\beta_{t}^{\prime}\mathbf{I})$ (13)
@@
$\mu_{t}^{\prime}(\bm{x}^{F}_{t},\bm{x}^{F}_{0})$ $=\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\bm{x}^{F}_{0}+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\bm{x}^{F}_{t}$ (14)
@@
To generate a sample from $q(\bm{x}^{F}_{0})$, we may sample from $q(\bm{x}^{F}_{T})$ and iteratively apply the reverse transitions $q(\bm{x}^{F}_{t-1}|\bm{x}^{F}_{t})$ back to $\bm{x}^{F}_{0}$. Given a sufficiently extensive $T$ and a suitably chosen $\beta_{t}$ series ($\beta_{t}\rightarrow 0$), the distribution $q(\bm{x}^{F}_{T})$ converges closely to an isotropic Gaussian. Specifically, our conditional diffusion model is parametrized through $p_{\theta}(x^{F}_{t-1}|x^{F}_{t},t,\tilde{x}^{F}_{0})$. Because $p_{\theta}(x^{F}_{t-1}|x^{F}_{t},t,\tilde{x}^{F}_{0})$ is not be easily estimated but could be approximated using a neural network ( i.e. DifFUSER), where the aim is to predict $\mu_{\theta}$ and $\Sigma_{\theta}$:
$p_{\theta}(\bm{x}^{F}_{t-1}|x^{F}_{t},t,\tilde{x}^{F}_{0})$ $\coloneqq\mathcal{N}(\bm{x}^{F}_{t-1};\mu_{\theta}(x^{F}_{t},t,\tilde{x}^{F}_{0}),\Sigma_{\theta}(x^{F}_{t},t,\tilde{x}^{F}_{0}))$ (16)
Here, a neural network can also be optimized to predict either $\epsilon$ or $\bm{x}^{F}_{0}$ [37, 10]. We opt for predicting $\bm{x}^{F}_{0}$ in this paper.
@@ 8.2 Pseudo-code of DifFUSER training and sampling
    $x^{F}_{0}\leftarrow$ concatenate(cam_feats, lidar_feats)
    diffusion_target $\leftarrow x^{F}_{0}$
@@
    $\tilde{x}^{F}_{0}\leftarrow\text{PDST}(x^{F}_{0})$
@@
    $x^{F}_{t}$leftarrow sqrt(alpha_cumprod(t) * $x^{F}_{0}$ + sqrt(1-alpha_cumprod(t)) * noise
@@
    $\hat{x}^{F}_{0}$leftarrow DifFUSER$(\tilde{x}^{F}_{0},t)$
@@
    pred $\leftarrow$ task_head($\hat{x}^{F}_{0}$)
@@
    loss $\leftarrow$ diffusion_loss(diffusion_target, $\hat{x}^{F}_{0}$)
@@
    $x^{F}_{0}\leftarrow$ concatenate(cam_feats, lidar_feats)
@@
        $\hat{x}^{F}_{0}\leftarrow$ DifFUSER($x^{F}_{0},x^{F}_{t},$ t_now)
    # estimate $x^{F}_{t}$ at t_next, sampler can be [37], [52], or [26]
        $x^{F}_{t}\leftarrow$ sampler_step($\hat{x}^{F}_{0}$, t_now, t_next)
@@
    pred $\leftarrow$ task_head($\hat{x}^{F}_{0}$)